How do u solve this question?
A study monitored the life spans of 129 robins over 4 years. The equation was developed to calculate the number of years y for x percent of the robin population to die.

y= 2-log(100-x)/.42
How many robins had died after 6 months?

Answer by ankor@dixie-net.com(12692)

A study monitored the life spans of 129 robins over 4 years. The equation was developed to calculate the number of years y for x percent of the robin population to die.
y= 2-log(100-x)/.42
How many robins had died after 6 months?
:
6 months = .5 years
:
(2 - log(100-x))/.42 = .5
:
2 - log(100-x) = .5 * .42; multiplied both sides by .42
:
2 - log(100-x) = .21
:
- log(100-x) = .21 - 2
:
- log(100-x) = -1.79
:
log(100-x) = 1.79; multiplied equation by -1
:
Find the anti-log (10^x) of both sides:
100 - x = 61.66
:
-x = 61.66 - 100
:
-x = -38.34; (*-1)
:
x = +38.34%; Therefore, .3834 * 129 = 49 birdies bit the dust in 6 mos
:
:
Check:
y = (2-log(100-38.34))/.42
:
y = (2 - log(61.66))/.42
:
y = (2 - 1.79)/.42
:
y = .4999 ~ .5